the distance diagonally across a square sandpit is 1.82 meters long . find the length of the sides of the sandpit

Answer by ewatrrr(24785)

(1.82m)^2 = 3.3124m^2
s^2 + s^2 = 3.3124
2s^2 = 3.3124 =
s^2 = 1.6562 (tossing out negative solution for unit measure)
s = 1.2869m